Let {an}∞n=1 be a convergent sequence with limit L such that an ≤ B forevery n≥N for some N ∈ N and some B ∈ R. Prove that L ≤ B


1
Expert's answer
2021-02-02T04:59:31-0500

We have that "\lim \limits_{n \to \infty} a_n = L"

Also, "a_n \leq B"

Taking limits we have "\lim \limits_{n \to \infty} a_n \leq \lim \limits_{n \to \infty} B"

Hence "L \leq \lim \limits_{n \to \infty} B = B"

As desired